OpenStudy (anonymous):

(-3,4) (-2,1) a. -5 b. -3 c. -3/5 d. -1 explain please

OpenStudy (anonymous):

you mean -3/5

OpenStudy (anonymous):

ok is there another step because I dont have that as an answer choice

OpenStudy (anonymous):

@satellite73 help

OpenStudy (anonymous):

you looking for the slope between the two poins \((-3,4) (-2,1) \)?

OpenStudy (anonymous):

form(-3\) to \(-2\) is one unit to the right on the \(x\) axis from \(4\) to \(1\) is \(3\) units down over 1, down 3, slope is \(-3\)

OpenStudy (anonymous):

@satellite73 yes i have to find its \[\frac{ rise }{ run }\]

OpenStudy (anonymous):

or you can compute \[\frac{4-1}{-3-(-2)}=\frac{3}{-1}=-3\] either way

OpenStudy (anonymous):

"rise" is \(-3\) and "run" is \(1\) so \(\frac{rise}{run}=-3\)

OpenStudy (anonymous):

@satellite73 oh ok thanks

OpenStudy (anonymous):

yw